OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.R(4, 0), m = 5

OpenStudy (anonymous):



OpenStudy (anonymous):

Use equation (y-y1)= m (x-x1) where (x1,y1) are the given co-ordinates & m is the slope of the line whose equation is to be find out. Here x1=4 & y1=0 & m=5 so putting the values in above equation we get y=5(x-4) (Answer)